14 men took 6 hours and 40 minutes to complete a job. if two of 8 men leave after working for one hour, how long will it take the remaining 6 men working at the same rate to complete the job?
Answer by ankor@dixie-net.com(22740)

14 men took 6 hours and 40 minutes to complete a job.
Find the number of man-hours required to do the job.
Change 6 hr 40 min to 6.6666 hrs
14 * 6.67 = 93.33 man-hrs require to complete the job
:
if two of 8 men leave after working for one hour, how long will it take the remaining 6 men working at the same rate to complete the job?
:
let h = time required by the remaining 6 men to complete the job
2(1) + 6(h+1) = 93.33
6h + 6 = 93.33 - 2
6h = 91.33 - 6
6h = 85.93
h = 85.93/6
h = 14.22 more hrs to completed the job
:
See if that checks out: 6(14.22+ 1) + 2(1) = 93.32 close enough